OpenStudy (anonymous):

A line has a slope of –3 and passes through the point (4, –3). Which is the equation of the line? Question 2 options: y = –3x – 15 y = –3x – 9 y = –3x + 9 y = –3x + 15

OpenStudy (anonymous):

Use the formula :(y-y1)=m(x-x1) here y1=-3 x1=4 and m=-3=slope

OpenStudy (anonymous):

im confused?

OpenStudy (dustinrathke24):

Plug your values into the formula. y-(-3)=-3(x-4)

OpenStudy (anonymous):

So its B?

OpenStudy (dustinrathke24):

So you have y+3=-3x+12 Move the 3 over so you have y=-3x+9

OpenStudy (anonymous):

Oh okay!

OpenStudy (dayakar):

equation of a line whose slope m=-3 and passing through the point(x1,y1) =(4,-3) is

OpenStudy (anonymous):

Thank you!

OpenStudy (dustinrathke24):

No problem.